OpenStudy (anonymous):

would this be right.. Ix-12I=3 15,-9 ??

OpenStudy (anonymous):

|x-12|=3 now we have two cases: x-12=3 or x-12=-3 solving them we get: x=3+12 x=-3+12 x=15 x=9

OpenStudy (anonymous):

could you help me with this one 

OpenStudy (anonymous):

is it 5, -7/3

OpenStudy (anonymous):

|x+6|=2x+1 x+6=2x+1 x+6=-(2x+1) x-2x=1-6 x+6=-2x-1 -x=-5 x+2x=-1-6 x=5 3x=-7 x=-7/3 Yes your result is correct. Well done :)

OpenStudy (anonymous):

sweet, thank you!

OpenStudy (anonymous):

My pleasure xD